# Efficient Decoder for End-to-End Oriented Object Detection in Remote Sensing Images

### v1

@@ III-C Selective Distinct Queries
where $i=1,2,\cdot\cdot\cdot,d$ and $d$ is the total number of decoder layers. After passing through all decoder layers, the model would typically have generated refined queries $q^{d}$. The loss $\mathcal{L}=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{L_{1}}\mathcal{L}_{1}+\lambda_{iou}\mathcal{L}_{iou}$ is applied at the queries of all the decoder layers for training, where $\mathcal{L}_{cls}$ is the focal loss [35] for classification, $\mathcal{L}_{1}$ and $\mathcal{L}_{iou}$ are L1 Loss and Rotated IoU/GIoU Loss [36, 37] for regression.
TABLE I: mAP results in different decoder layers. Sparse R-CNN (100 queries) without random resize data processing are trained on COCO train set and per layer mAP is reported on validation set.
@@
Sparse R-CNN 14.1 31.6 39.7 43.7 44.7 45.0
Selective Queries. As shown in Table I, it appears that there is a clear growing trend in mAP as queries move deeper into the decoder layers. Notably, the mAP for the last few layers is very close to each other. Take Sparse R-CNN as an example, the mAP is 43.7%, 44.7% and 45.0% at the $4^{th}$, $5^{th}$ and $6^{th}$ decoder layer respectively. This suggests that the queries in the last few decoder layers already capture sufficient information about object positions and categories. Therefore, we can collect the queries from $p$ intermediate decoder layers as selective queries only once and similar queries among them will be filtered to obtain distinct queries in the next step, shown in Fig. 3(b).
@@
Distinct Queries. The selective queries collected from intermediate decoder layers already capture sufficient information about the positions and categories with the images. But there exist many similar queries. To obtain distinct queries, similar queries will be filtered,
@@
where $\hat{q}$ denotes distinct queries. Distinct queries are calculated in two steps. First, we filter similar queries by a simple class-agnostic non-maximum suppression (NMS), denoted as $\psi(\cdot)$. Second, if the remaining queries $\psi(q^{p})$ are more than the initial queries $n$, we will choose the first $n$ ones according to scores as distinct queries. Otherwise, all remaining queries are treated as distinct queries. The class-agnostic NMS operation here is regarded as the first step of distinct queries mechanisms embedded among intermediate decoder layers, instead of as post-processing at the inference stage in traditional detectors [3, 4]. Therefore, such a method still conforms to the end-to-end standard. In the basic queries update way, the last decoder layer $D^{d}$ exclusively refines queries from the $D^{d-1}$, while our method handles queries from multiple intermediate layers simultaneously. Through filtering similar queries, the number of distinct queries is reduced which reduces computational overhead.
@@
Figure 4: The ratios of similar queries under different high IoU threshold. Compared with baseline, SDQ significantly remove similar queries.
@@
Analysis of Selective Distinct Queries. Our proposed SDQ can filter similar queries and keep the remaining queries distinct. The one-to-one label assignment, (e.g. Hungarian matching) was first introduced to object detection by DETR [6]. Only one sample can be assigned to a corresponding ground truth object according to matching cost matrices, while all other samples are assigned to background class. However, there exists a large number of similar queries which hinder training optimization. We use Intersection over Union (IoU) to measure the similarity of queries at the stage of one-to-one label assignment. If the IoU of two predict boxes is greater than the threshold, their corresponding queries are regarded as similar. Why only use IoU to measure the similarity of queries, rather than considering all matching cost matrices including IoU, L1 distance and Focal loss? There are two main reasons. First, the model adopts a one-to-one label assignment. Second, there are no ground truth objects with an IoU greater than a threshold close to 1, e.g. 0% ground truth objects with an IoU greater than 0.8 in the DIOR-R test set [38]. If a positive sample is assigned to an object, any other query with an IoU greater than the threshold is assigned as a background class. To quantify the similarity among queries, we conduct an analytical experiment on the DIOR-R test set shown in Fig. 4. Take threshold = 0.95 as an example, 4.00% of queries showed similarity in baseline model DDQ-O [12]. Our SDQ can significantly reduce the value to 0.74% which can prove the effectiveness of our method.
@@
TABLE II: Comparison with state-of-the-art methods on the DIOR-R dataset.
@@
RetinaNet-O [35] 61.49 28.52 73.57 81.17 23.98 72.54 19.94 72.39 58.20 69.25 79.54 32.14 44.87 77.71 67.57 61.09 81.46 47.33 38.01 60.24 57.55
Oriented Rep [39] 70.03 46.11 76.12 87.19 39.14 78.76 34.57 71.80 80.42 76.16 79.41 45.48 54.90 87.82 77.03 68.07 81.60 56.83 51.57 71.25 66.71
DCFL [1] 68.60 53.10 76.70 87.10 42.10 78.60 34.50 71.50 80.80 79.70 79.50 47.30 57.40 85.20 64.60 66.40 81.50 58.90 50.90 70.90 66.80
@@
Faster RCNN-O [27] 62.79 26.80 71.72 80.91 34.20 72.57 18.95 66.45 65.75 66.63 79.24 34.95 48.79 81.14 64.34 71.21 81.44 47.31 50.46 65.21 59.54
Gliding Vertex [40] 65.35 28.87 74.96 81.33 33.88 74.31 19.58 70.72 64.70 72.30 78.68 37.22 49.64 80.22 69.26 61.13 81.49 44.76 47.71 65.04 60.06
RoI Transformer [34] 63.34 37.88 71.78 87.53 40.68 72.60 26.86 78.71 68.09 68.96 82.74 47.71 55.61 81.21 78.23 70.26 81.61 54.86 43.27 65.52 63.87
QPDet [41] 63.22 41.39 71.97 88.55 41.23 72.63 28.82 78.90 69.00 70.07 83.01 47.83 55.54 81.23 72.15 62.66 89.05 58.09 43.38 65.36 64.20
AOPG [38] 62.39 37.79 71.62 87.63 40.90 72.47 31.08 65.42 77.99 73.20 81.94 42.32 54.45 81.17 72.69 71.31 81.49 60.04 52.38 69.99 64.41
@@
RRoIFormer 67.31 55.23 74.19 82.74 44.49 78.56 39.85 70.27 79.84 75.10 80.38 45.64 58.51 88.91 68.10 75.73 85.52 57.17 53.54 65.05 67.31
TABLE III: Comparison with state-of-the-art methods on the DOTA-v1.0 dataset. * denotes multi-scale trainig and testing. D. DETR represents Deformable DETR.
@@
Mask OBB R50 89.61 85.09 51.85 72.90 75.28 73.23 85.57 90.37 82.08 85.05 55.73 68.39 71.61 69.87 66.33 74.86
SCRDet [47] R101 89.98 80.65 52.09 68.36 68.36 60.32 72.41 90.85 87.94 86.86 65.02 66.68 66.25 68.24 65.21 72.61
RoI Transformer [34] R50 88.65 82.60 52.53 70.87 77.93 76.67 86.87 90.71 83.83 52.81 53.95 67.61 74.67 68.75 61.03 74.61
Gliding Vertex [40] R101 89.64 85.00 52.26 77.34 73.01 73.14 86.82 90.74 79.02 86.81 59.55 70.91 72.94 70.86 57.32 75.02
Oriented R-CNN [3] R50 89.46 82.12 54.78 70.86 78.93 83.00 88.20 90.90 87.50 84.68 63.97 67.69 74.94 68.84 52.28 75.87
@@
D. DETR-O [7] R50 78.95 68.64 32.57 55.17 72.53 57.77 73.71 88.36 75.46 79.34 45.36 53.78 52.94 66.35 50.38 63.42
D. DETR-O w/ CSL [48] R50 86.27 76.66 46.64 65.29 76.80 76.32 87.74 90.77 79.38 82.36 54.00 61.47 66.05 70.46 61.97 72.15
ARS-DETR [42] R50 86.61 77.26 48.84 66.76 78.38 78.96 87.40 90.61 82.76 82.19 54.02 62.61 72.64 72.80 64.96 73.79

### v4

@@ 3.3 Selective Distinct Queries
where $i=1,2,\cdot\cdot\cdot,d$ and $d$ is the total number of decoder layers. After passing through all decoder layers, the model would typically have generated refined queries $q^{d}$. The loss $\mathcal{L}=\lambda_{cls}\mathcal{L}_{cls}+\lambda_{L_{1}}\mathcal{L}_{1}+\lambda_{iou}\mathcal{L}_{iou}$ is applied at the queries of all the decoder layers for training, where $\mathcal{L}_{cls}$ is the focal loss (Lin et al., 2017) for classification, $\mathcal{L}_{1}$ and $\mathcal{L}_{iou}$ are L1 Loss and Rotated IoU (Zhou et al., 2019) for regression.
Table 1: mAP results in different decoder layers. ARS-DETR and Deformable DETR-O with 300 queries are trained on DOTA-v1.0 training and validation set, and the mAP at every decoder layer is reported on test set.
@@
ARS-DETR 0.40 3.47 19.56 33.07 39.63 72.66
Deformable DETR-O 0.13 5.49 23.64 30.00 36.61 63.42
Selective Queries. As shown in Table 1, it appears that there is a clear growing trend in mAP as queries move deeper into the decoder layers. Notably, the mAP can be easily tested by adjusting the number of decoder layers, since the outputs of each decoder layer are supervised by loss functions. Taking ARS-DETR (Zeng et al., 2024) as an example, the mAP is 33.07% and 39.63% at the $4^{th}$ and $5^{th}$ decoder layer respectively. This suggests that the queries in the last few decoder layers already capture sufficient information about object positions and categories. Therefore, we can collect the queries from $p$ intermediate decoder layers as selective queries only once, and similar queries among them will be filtered to obtain distinct queries in the next step, shown in Fig. 4(b).
@@
Distinct Queries. The selective queries collected from intermediate decoder layers already capture sufficient information about the positions and categories with the images. But there exist many similar queries. To obtain distinct queries, similar queries will be filtered according to:
@@
where $\hat{q}$ denotes distinct queries. Distinct queries are calculated in two steps. First, we filter similar queries by a simple class-agnostic non-maximum suppression (NMS), denoted as $\psi(\cdot)$. Second, if the remaining queries $\psi(q^{p})$ are more than the initial queries $n$, we will choose the first $n$ ones according to scores as distinct queries. Otherwise, all remaining queries are treated as distinct queries. The class-agnostic NMS operation here is regarded as the first step of distinct queries mechanisms embedded among intermediate decoder layers, instead of as post-processing at the inference stage in traditional detectors (Xie et al., 2021; Han et al., 2021b). Therefore, such a method still conforms to the end-to-end standard. In the basic queries update way, the last decoder layer $D^{d}$ exclusively refines queries from the $D^{d-1}$, while our method handles queries from multiple intermediate layers simultaneously. Through filtering similar queries, the number of distinct queries is reduced which reduces computational overhead.
@@
Figure 5: The ratios of similar queries under different high IoU threshold on DIOR-R. Compared with baseline, SDQ significantly filters similar queries.
@@
Analysis of Selective Distinct Queries. Our proposed SDQ can filter similar queries and keep the remaining queries distinct. The one-to-one label assignment was first introduced to object detection by DETR (Carion et al., 2020). Only one sample can be assigned to a corresponding ground truth object according to matching cost matrices, while all other samples are assigned to the background class. However, there exists a large number of similar queries which hinder training optimization. We use Intersection over Union (IoU) to measure the similarity of queries at the stage of one-to-one label assignment. If the IoU between a positive sample and a background query is greater than the threshold, they are regarded as similar. We define the ratios of similar queries as:
$\mathrm{ratios}=\frac{\sum_{i=1}^{N_{\mathrm{pos}}}\mathbbm{1}(\max(\mathrm{IoU}(\mathrm{pos}_{i},\mathrm{neg}_{j}))>t)}{\sum_{i=1}^{N_{\mathrm{GT}}}\mathrm{GT}_{i}}$ (9)
where $\mathbbm{1}(\cdot)$, $\mathrm{pos}$, $\mathrm{neg}$, $t$, $\mathrm{GT}$ denote indicator function, positive samples, negative samples, threshold, and ground truth objects. Note that the number of positive samples $N_{\mathrm{pos}}$ and ground truth objects $N_{\mathrm{GT}}$ is always equal. Why only use IoU to measure the similarity of queries, rather than considering all matching cost matrices including IoU, L1 distance, and Focal loss? There are two main reasons. First, the model adopts the one-to-one label assignment. Second, there are no ground truth objects with an IoU greater than a threshold close to 1, e.g., 0% ground truth objects with an IoU greater than 0.8 in the DIOR-R test set (Cheng et al., 2022). To quantify the similarity among queries, we conduct an analytical experiment on the DIOR-R test set shown in Fig. 5. Taking $t=0.95$ as an example, 4.00% of queries showed similarity in baseline model DDQ-O (Zhang et al., 2023b). Our SDQ can significantly reduce the value to 0.74% which can prove the effectiveness of our method.
3.4 Prediction, Label Assignment and Loss
In each decoder layer, the model predicts categories and boxes. This prediction is achieved through a simple Feedforward Neural Network (FFN) and linear layers.
Object detection tackles two sub-tasks, regression for positions and classification for categories. To execute these tasks, bipartite matching is applied at the label assignment stage between predictions and ground truth objects. The matching cost is calculated as follows:
$\mathcal{L}=\lambda_{iou}\cdot\mathcal{L}_{iou}+\lambda_{L_{1}}\cdot\mathcal{L}_{L_{1}}+\lambda_{cls}\cdot\mathcal{L}_{cls}$ (10)
where $\mathcal{L}_{iou}$ and $\mathcal{L}_{L_{1}}$ are Rotate IoU loss (Zhou et al., 2019) and $L_{1}$ loss between predicted boxes and ground truth boxes, respectively. $\mathcal{L}_{cls}$ is focal loss (Lin et al., 2017) for classification between predicted categories and ground truth labels. $\lambda_{iou}$, $\lambda_{L_{1}}$ and $\lambda_{cls}$ are coefficient of corresponding component. Based on matching cost $\mathcal{L}$, only one optimal sample is assigned to each ground truth object, while the remaining samples are assigned to the background.
Training losses are mathematically the same as matching costs, but they are performed on the matched pairs. The losses are applied to all predicted content and positional queries across all decoder layers.
4 EXPERIMENTS
4.1 Datasets
We conduct our experiments on six public datasets, including oriented object detection datasets DIOR-R (Cheng et al., 2022), DOTA-v1.0 (Xia et al., 2018)/v1.5/v2.0 (Ding et al., 2021), a scene text detection dataset ICDAR2015 (Karatzas et al., 2015), and a horizontal object detection dataset MS COCO (Lin et al., 2014).
DIOR-R (Cheng et al., 2022) is a large-scale oriented object detection dataset. It contains a total of 23,463 remote sensing images and 192,512 instances covering 20 common classes including windmill (WM), trainstation (TS), storagetank (STO), ship (SH), harbor (HA), golffield (GF), expressway toll station (ETS), bridge (BR), baseballfield (BF), airplane (APL), vehicle (VE), tenniscourt (TC), stadium (STA), overpass (OP), groundtrackfield (GTF), dam (DAM), expressway service area (ESA), chimney (CH), basketballcourt (BC) and airport (APO). We train our model on training and validation sets and report the results on the test set.
DOTA series including DOTA-v1.0/v1.5/v2.0 are used. DOTA-v1.0 (Xia et al., 2018) is designed for oriented object detection tasks which contain 188,282 instances and 2,806 remote sensing images ranging in size from 800$\times$800 to 4,000$\times$4,000. It covers 15 different object categories: Plane (PL), Baseball diamond (BD), Bridge (BR), Ground track field (GTF), Small vehicle (SV), Large vehicle (LV), Ship (SH), Tennis court (TC), Basketball court (BC), Storage tank (ST), Soccer-ball field (SBF), Roundabout (RA), Harbor (HA), Swimming pool (SP), and Helicopter (HC). DOTA-v1.5 uses the same images as DOTA-v1.0 but adds a new category Container Crane (CC) and more small instances, which contains 403,318 instances. DOTA-v2.0 adds two categories Airport and Helipad compared to DOTA-v1.5, which contains 11,268 images and 1,793,658 instances. We use both training and validation sets for training and test sets for testing for the DOTA series. The detection result of the test set is obtained by submitting testing results to DOTA’s official evaluation server.
ICDAR2015 (Karatzas et al., 2015) is a text detection dataset, which includes 1,000 training images and 500 test images.
MS COCO 2017 (Lin et al., 2014) is a horizontal object detection dataset for natural scenes, consisting of 118K images as train2017 for training and 5K images as val2017 for validation. We train our model on train2017 and report the evaluation results on val2017.
4.2 Implementation Details
We select ResNet-50 (He et al., 2016) as the backbone which is pre-trained on the ImageNet (Deng et al., 2009). We optimize the overall network with AdamW optimizer (Loshchilov and Hutter, 2018) with the warming up for 500 iterations. We use Focal loss (Lin et al., 2017) for classification and Rotated IoU/GIoU loss (Zhou et al., 2019) and L1 loss for regression.
For experiments on the DOTA-v1.5/v2.0 and ICDAR2015, we use 2 NVIDIA RTX 2080ti with a batch size of 4 (2 images per GPU) and the learning rate of $5\times e^{-5}$ for training, while a single GPU with a batch size of 2 and the learning rate of $2.5\times e^{-5}$ on DOTA-v1.0 and DIOR-R. Models build on MMRotate (Zhou et al., 2022b) with Pytorch. The weights of losses are 2.0, 2.0, and 5.0 for Focal loss, L1 loss, and Rotated IoU loss respectively. Following the official settings of the DOTA benchmark, we crop images into patches of $1024\times 1024$ with overlaps of 200 and train the model for 24 epochs. Images on DIOR-R are trained for 36 epochs with the original fixed size of $800\times 800$, while 12 epochs in ablation studies. To compare with other methods, the number of queries is 500, while in ablation studies on DIOR-R, it is 300. Images on ICDAR2015 are trained for 160 epochs with a fixed size of $800\times 800$.
For experiments on COCO 2017, we use 2 NVIDIA RTX 2080ti with a batch size of 4 (2 images per GPU) for training. Models build on MMDetection (Zhou et al., 2022b) with Pytorch. The weights of loss are 2.0, 5.0, and 2.0 for Focal loss, L1 loss, and IoU loss respectively. The learning rate is $2\times e^{-5}$.
Table 2: RQFormer ablation experiments with ResNet-50 on DIOR-R. The best result in each group is colored gray.
Pooling Size 3 5 7 9
mAP 61.02 61.35 62.63 62.32
Params(M) 40.71 40.78 40.88 41.01
(a) Effect of pooling size of RRoI Attention. Gradually increasing the pooling size, the performance is the best when it is 7.
Head 2 4 8 16
mAP 61.36 62.19 62.63 55.63
Params(M) 40.73 40.78 40.88 41.08
(b) Effect of heads of RRoI Attention. Gradually increasing the number of heads, the performance is saturated when it is 8.
Attention mAP Params(M) FPS
Instance interact 62.41 63.68 14.12
Dynamic Attention 61.59 46.97 11.70
RRoI Attention 62.76 40.88 14.28
(c) RRoI attention vs. other cross-attention. Our RRoI attention outperforms dynamic attention and instance interaction.
threshold 0.8 0.85 0.9 0.95
mAP 62.58 62.76 62.63 62.49
Params(M) 40.88 40.88 40.88 40.88
(d) Effect of IoU threshold of SDQ. Increasing the threshold, the best performance occurs when it is 0.85.
Method Layers Selective Distinct mAP
Top-$n$ 3 $1^{st}$, $2^{nd}$ $2^{nd}$ 59.49
Stack 3 $1^{st}$, $2^{nd}$ $2^{nd}$ 59.99
SDQ 3 $1^{st}$, $2^{nd}$ $2^{nd}$ 60.66
(e) SDQ vs. other query selection methods. Our proposed SDQ outperforms stack and top-$n$ method.
Method Layers Selective Distinct mAP
SDQ 3 $2^{nd}$ $2^{nd}$ 62.08
SDQ 2 $1^{st}$ $1^{st}$ 62.63
SDQ 1 $0^{th}$ $0^{th}$ 44.02
(f) Effects of number of decoder layers. Gradually increasing the number of layers, the performance is saturated at 2 layers.
Method Layers Selective Distinct mAP Params(M)
SDQ 3 $0^{th}$, $1^{st}$, $2^{nd}$ $2^{nd}$ 58.01 42.70
SDQ 3 $1^{st}$, $2^{nd}$ $2^{nd}$ 60.66 42.70
SDQ 3 $1^{st}$ $1^{st}$ 58.03 42.70
SDQ 3 $2^{nd}$ $2^{nd}$ 62.08 42.70
(g) Different combinations of selective queries and distinct queries in the SDQ.
Method RRoI Att. SDQ mAP Params(M) FPS
DDQ-O(baseline) 61.66 63.68 14.40
RQFormer 62.31 40.88 14.76
RQFormer 62.41 63.68 14.12
RQFormer 62.76 40.88 14.28
(h) Ablation study on each components in RQFormer. Starting from DDQ-O, we gradually add RRoI attention and SDQ.
4.3 Ablation Study
Different pooling size. To further explore RRoI Attention, we adopt different pooling sizes $r$. As the pooling size $r$ increases, more features will be highlighted in positional queries. When pooling size $r$ is set to 5, the AP is only 61.35% on DIOR-R, shown in Table 2 (a). When it is increased to 7, the mAP can achieve 62.63%. This suggests that abundant features are sufficient to represent objects. When we increase the pooling size to 9, the performance of the model declines. It indicates that redundant features confuse the attention mechanism. As the pooling size gradually increases, the model’s parameters also increase, given that the parameters of values $V$ become larger.
The number of attention head. Following deformable attention (Zhu et al., 2020), our RRoI attention also adopts multiple heads. Different heads establish different associations between queries and values. In this ablation, we use different numbers of $m$ heads. As we increase the number of heads from 2 to 8, the mAP grows from 61.36% to 62.63% on DIOR-R, shown in Table 2 (b). This indicates that the multi-head attention extends the ability to focus on different parts of features, and provides multiple subspaces for representation. When we increase the number of heads to 16, there is a decline in performance. We suspect that the redundant head focuses on redundant features. As the number of heads increases, the model’s parameter count rises, driven by the increased number of parameters in values $V$, attention weights $A$, and projection matrices.
The RRoI attention vs. other cross-attention. To further affirm the efficacy of RRoI attention, we conduct a comparative analysis with other attention mechanisms, namely instance interaction (Sun et al., 2021) and dynamic attention (Dai et al., 2021), shown in Table 2 (c). Although instance interaction is not strictly an attention mechanism, we discuss it for brevity. RRoI attention attains an mAP of 62.76%, outperforming instance interaction by 0.35% and dynamic attention by 1.17%. The model incorporating RRoI attention exhibits the lowest parameter count, standing at 40.88M, which is 22.8M less than instance interaction and 6.09M less than dynamic attention. This reduction is attributed to RRoI attention’s focus on the region inside boxes, resulting in significant computational efficiency. Moreover, the model with RRoI attention achieves faster detection speed, reaching 14.28 FPS, surpassing both instance interaction and dynamic attention.
Threshold of Selective Distinct Queries. To further explore the effectiveness of SDQ, we conducted ablation experiments on the different IoU thresholds shown in Table 2 (d). To filter the similar queries and keep the remaining queries distinct, the threshold is a high value close to 1 in SDQ. As the IoU threshold value increases from 0.8 to 0.85, the mAP increases by 0.18% on DIOR-R. When the IoU threshold increases to 0.95 which is very close to 1, the mAP value decreases slightly, indicating that there are very few predicted boxes that fully coincide or are close to perfectly coinciding. The number of parameters is the same at different thresholds.
@@ 4.3 Ablation Study
Table 3: Comparison with state-of-the-art methods on the DIOR-R. The results in bold denote the best performance of each column.
@@
RetinaNet-O 61.49 28.52 73.57 81.17 23.98 72.54 19.94 72.39 58.20 69.25 79.54 32.14 44.87 77.71 67.57 61.09 81.46 47.33 38.01 60.24 57.55
DFDet 61.92 38.83 77.41 81.36 34.11 74.94 26.26 62.31 76.06 75.56 79.62 38.26 52.76 80.40 73.11 68.27 81.38 52.23 44.11 63.35 62.11
SOOD 62.61 46.82 72.00 81.28 41.40 72.56 78.63 68.91 27.53 78.64 81.35 46.88 56.32 80.57 73.74 68.36 81.43 56.91 47.33 65.52 64.41
Oriented Rep 70.03 46.11 76.12 87.19 39.14 78.76 34.57 71.80 80.42 76.16 79.41 45.48 54.90 87.82 77.03 68.07 81.60 56.83 51.57 71.25 66.71
DCFL 68.60 53.10 76.70 87.10 42.10 78.60 34.50 71.50 80.80 79.70 79.50 47.30 57.40 85.20 64.60 66.40 81.50 58.90 50.90 70.90 66.80
@@
Faster RCNN-O 62.79 26.80 71.72 80.91 34.20 72.57 18.95 66.45 65.75 66.63 79.24 34.95 48.79 81.14 64.34 71.21 81.44 47.31 50.46 65.21 59.54
Gliding Vertex 65.35 28.87 74.96 81.33 33.88 74.31 19.58 70.72 64.70 72.30 78.68 37.22 49.64 80.22 69.26 61.13 81.49 44.76 47.71 65.04 60.06
RoI Transformer 63.34 37.88 71.78 87.53 40.68 72.60 26.86 78.71 68.09 68.96 82.74 47.71 55.61 81.21 78.23 70.26 81.61 54.86 43.27 65.52 63.87
QPDet 63.22 41.39 71.97 88.55 41.23 72.63 28.82 78.90 69.00 70.07 83.01 47.83 55.54 81.23 72.15 62.66 89.05 58.09 43.38 65.36 64.20
AOPG 62.39 37.79 71.62 87.63 40.90 72.47 31.08 65.42 77.99 73.20 81.94 42.32 54.45 81.17 72.69 71.31 81.49 60.04 52.38 69.99 64.41
@@
DDQ-O (baseline) 66.71 54.14 73.11 81.80 45.25 77.60 33.37 70.09 79.18 72.24 75.42 47.18 58.27 89.50 70.98 74.62 82.83 55.03 56.23 66.61 66.51
RQFormer 67.31 55.23 74.19 82.74 44.49 78.56 39.85 70.27 79.84 75.10 80.38 45.64 58.51 88.91 68.10 75.73 85.52 57.17 53.54 65.05 67.31
Table 4: Comparison with state-of-the-art methods on the DOTA-v1.0 dataset. * denotes multi-scale training and testing. D. DETR represents Deformable DETR. The results in bold denote the best performance of each column.
@@
H2RBox R50 88.24 79.30 42.76 55.79 78.90 72.70 77.54 90.85 81.96 84.38 55.28 64.49 61.91 70.63 51.51 70.41
SCRDet R101 89.98 80.65 52.09 68.36 68.36 60.32 72.41 90.85 87.94 86.86 65.02 66.68 66.25 68.24 65.21 72.61
RoI Transformer R50 88.65 82.60 52.53 70.87 77.93 76.67 86.87 90.71 83.83 52.81 53.95 67.61 74.67 68.75 61.03 74.61
Gliding Vertex R101 89.64 85.00 52.26 77.34 73.01 73.14 86.82 90.74 79.02 86.81 59.55 70.91 72.94 70.86 57.32 75.02
Oriented R-CNN R50 89.46 82.12 54.78 70.86 78.93 83.00 88.20 90.90 87.50 84.68 63.97 67.69 74.94 68.84 52.28 75.87
@@
D. DETR-O R50 78.95 68.64 32.57 55.17 72.53 57.77 73.71 88.36 75.46 79.34 45.36 53.78 52.94 66.35 50.38 63.42
D. DETR-O w/ CSL R50 86.27 76.66 46.64 65.29 76.80 76.32 87.74 90.77 79.38 82.36 54.00 61.47 66.05 70.46 61.97 72.15
EMO2-DETR R50 88.08 77.91 43.17 62.91 74.01 75.09 79.21 90.88 81.50 84.04 51.92 59.44 64.74 71.81 58.96 70.91
ARS-DETR R50 86.61 77.26 48.84 66.76 78.38 78.96 87.40 90.61 82.76 82.19 54.02 62.61 72.64 72.80 64.96 73.79